OpenStudy (anonymous):

if given just the volume of a cylinder, how do you find the radius and the height so you can build a model???

OpenStudy (yuki):

you need more information that just the volume since the volume of a cylinder depends of the radius of the base and the height of the cylinder

OpenStudy (yuki):

if we know the relationship between the height and the radius, we can figure out what they are

OpenStudy (anonymous):

i have to make either one model that fits both a surface area and volume, or two, one that fits each the only things i have been given to go on a volume wich is 192 times pi(my teacher deoesn't use 3.14) and a total surface area wich is 112 times pi

OpenStudy (anonymous):

i have been working on this for hours and am soooooo frustrated

OpenStudy (anonymous):

i am so stuck, please help

OpenStudy (yuki):

all right so a volume of a cylinder is represented by \[\pi r^2 *h\] and the surface area is represented by \[2\pi r^2 +2 \pi r h\]

OpenStudy (yuki):

so we can have the following system of eqn.s \[\pi r^2h = 192 \pi\]

OpenStudy (yuki):

and \[2\pi r ^2 + 2\pi r h = 112 \pi\]

OpenStudy (yuki):

all you have to do now is solve the system from the first eqn. \[h = {192 \over r^2}\] if I substitute that in the second eqn. \[2\pi r^2 + 2\pi r * {192 \over r^2}= 112\]

OpenStudy (dumbcow):

V = pi*r^2*h = 192pi SA = 2pi*r^2 +2pi*r*h = 112pi canceling out the pi r^2*h = 192 2(r^2+rh) = 112 from 1st equation h=192/r^2 substitute and solve for r in 2nd equation r^2 +r(192/r^2) = 56 r^2 +192/r = 56 (r^3 +192)/r = 56 r^3-56r +192 = 0 working on seeing if this factors

OpenStudy (yuki):

I can simplify this into\[2\pi r^2 + 384\pi / r = 112\] so you will get \[r^3-56r+192 = 0\]

OpenStudy (anonymous):

thats pretty much where im at now, but translating that into something i can make a model from i have no idea how to do

OpenStudy (yuki):

what do you mean? once you solve for r, you just plug that in to the original eqn to find what h is

OpenStudy (anonymous):

i am not sure i am solving for r correctly then, the numbers i am comming up with just don't seem right

OpenStudy (dumbcow):

does not factor using a graph, got approximation r=8.82 so h=2.468

OpenStudy (yuki):

yea, I tried the rational root theorem but they all failed.

OpenStudy (anonymous):

this teacher suks!!

OpenStudy (dumbcow):

what type of model do you have to do, you could use the relationship between the r and h found to scale any similar cylinder of any size

OpenStudy (yuki):

if you want to find one that fits only one, then you let r=1 then h =192 if you let h =1 then r = sqrt(192)

OpenStudy (anonymous):

i can make it out of anything, just has to be either one cylinder that has a volume of 192times pi and one that has a surface area of 112 times pi, or one that equals both measurements

OpenStudy (yuki):

then the one I just recommended works

OpenStudy (yuki):

one that equals both measures is a tough one because r is not an exact number

OpenStudy (anonymous):

im still stmuped

OpenStudy (yuki):

what do you mean ?

OpenStudy (yuki):

just use r = 1 and h =192 and r = 1 and h = 55

OpenStudy (dumbcow):

i made a mistake too the only solution for r is -8.82 not positive so really there is no solution

OpenStudy (anonymous):

thank you yuki and cow, you were both a lot of help

OpenStudy (dumbcow):

go with yuki's idea and pick an h and r to satisfy one of the 2 conditions your welcome

OpenStudy (yuki):

good luck :) and when you are spending more than 2 hours on your homework, you are pushing yourself too far rest a little, ask your friends, your instructors. and don't forget that we are also here for you